# A Hamiltonian for 1:1 Rotational Secondary Resonances, and Application to Small Satellites of Saturn and Jupiter

### v1

@@ 3.3 Atlas of the phase space
Figure 4: The surfaces of section (SOS; first column), and level curves of the hamiltonian (29) (second column), and (33) (third column) in the cases of Aegaeon and Prometheus (top and bottom, respectively). The separatrix of the $\alpha$-regime is shown in red. Dashed-dot-horizontal lines indicate the fixed points obtained from hamiltonians and compared to the SOS corresponding values.
3.5 Additional application: Aegaeon and Prometheus.
@@
Fig. (4) shows the rotational phase space around the synchronous regime in the cases of Aegaeon ($\epsilon\sim 1.5236$), and Prometheus ($\epsilon\sim 1.1727$). As pointed out in Section 1, both rotational phase spaces display the $\alpha$ and $\beta$-regimes. Inspection of the left and middle panels shows us that the position of the three fixed points of the $\alpha$ and $\beta$-regimes obtained from level curves (middle) agree with surfaces of section in the case of Prometheus, but the same is not true in the case of Aegaeon [^5] Note that our model does not allow to obtain the loci of the separatrix of the synchronous regime since it is valid, by construction, only in the interior of the resonance..
@@
To improve the model we include terms of order higher than $J^{2}$ in the actions so that the hamiltonian (24) becomes:
@@
$-\frac{1}{128\epsilon}(x^{2}+y^{2})^{3}+\frac{17}{96\sqrt{\epsilon}}e(x^{2}+y^{2})^{5/2},$ (33)
where terms up to $J^{3}$ have been included (corresponding to the last two terms in Equation (33). A significant improvement can be seen, mainly in the case of Aaegeon. As discussed in Lei (2023), and first shown in Gkolias et al. (2019), in the specific case of the 1:1 secondary resonance, the analytical and numerical mapping of the associated fixed points diverge for values of $\epsilon>1.2$.
@@ 4 Conclusions
This work shows the results of the dynamics of rotation of out-of-round close-in small Saturnian satellite Methone and Aegaeon discovered by the Cassini spacecraft. Their shapes and physical parameters have been updated recently by Thomas and Helfenstein (2020). As pointed out in Section 1, the main problem we are considering in this work is related to the nature of the $\alpha$ and $\beta$-regimes of motion in the phase space of rotation of these satellites already detected for Amalthea, Metis, and Prometheus (and probably others - see Melnikov and Shevchenko 2022, and references therein).
@@
We first conduct our investigation with the well-known model of rotation developed by Goldreich and Peale (1966) and applied by many authors in distinct sorts of problems through decades (e.g. Wisdom et al. 1984, Wisdom 2004, Callegari and Rodriguez 2013). The exact equations of the 1.5 degree-of-freedom model are solved numerically Everhart (1985) code, given very accurate rotational trajectories able to identify the main properties of the phase space with the surface of section technique. The $\beta$-regime is located close to the origin and corresponds to the “classical” fixed point associated with the synchronous rotation motion where the amplitude of oscillation of the physical angle $\Psi=\theta-f$ is null at the equilibrium. In the case of Methone, its center is slightly forced by an amount $\sim 0.02$ such that $\frac{\dot{\theta}}{n}\sim 1.02$, a value which can be confirmed by linear theory (Wisdom et al. 1984; Callegari and Rodriguez 2013). The $\alpha$-regime engulfs the $\beta$-regime and it is located far from origin at $\frac{\dot{\theta}}{n}\sim 0.45$ such that $\frac{\dot{\theta}}{n}$ oscillates around 1. Also interestingly the amplitude of variation of $\Psi$ is not null in the equilibrium point of the $\alpha$-regime keeping an amplitude of the order of 30 degrees.
@@
To understand the complexity of such rotation phase space of Methone and Aegaeon, we apply the model developed by Wisdom (2004) to study the rotation of Enceladus. Thus, we follow exactly his steps and reproduce his results; having completed this task, we could apply the methodology to our case. Our hamiltonian (18) generalizes Wisdom’s theory showing all terms at that order of expansion (second order in action). Hamiltonian (33) includes also terms up to $J^{3}$ in the actions. The terms in the expansion involving those which are proportional to 1:1 resonance between the frequency of physical libration $\Psi$ and the mean motion $n$ are responsible for the co-existence of the $\alpha$ and $\beta$ regimes in the rotation of the current Methone and Aegaeon’s phase spaces. Therefore, we have shown the existence of a secondary resonance within the synchronism, completing our initial goals.
@@
Figures (2) and (3) show the dependence of the domains of the $\alpha$ and $\beta$ regimes with the shape parameter of the Methone. The $\beta$-regime occurs at $\epsilon=\epsilon_{c}\sim 1.0149$ such that for $\epsilon<\epsilon_{c}$ only the $\alpha$-regimes exists. While there is good agreement between numerical and analytical calculations in the case of Methone, the same is not true for Aegaeon, where the model is useful only when higher order terms are included in the hamiltonian since $epsilon>1.2$, a superior limit for convergence of analytical estimative (Gkolias et al. 2019).
@@
As we pointed out above, trajectories within this regime of motion are forced and suffer “long-term” oscillations even in the case of initial conditions located over the fixed point. Being that the current rotational states of Saturnian close-in small satellites are probably synchronous (Thomas and Helfenstein 2020), our results can be taken into account in evolutionary studies of the rotation of these bodies. Tidal models (e.g. Ferraz-Mello et al. 2008, Ferraz-Mello 2013) can be applied to estimate the final destiny of the rotation of such small bodies and also the role of the secondary resonances on its thermal emission (Wisdom 2004).
@@ ACKNOWLEDGEMENTS.
Prof. Adrián Rodríguez (Valongo Observatory, UFRJ/Brazil).
Part of this work has been presented at the ‘XXI Brazilian Colloquium on Orbital Dynamics’ (INPE, from December 12 to 16, 2022); and at the ‘New Frontiers of Celestial Mechanics: theory and applications’ (Department of Mathematics Tullio Levi-Civita, University of Padua, from February 15 to 17, 2023).

### v3

@@ 3.5 Amplitude of the physical libration
Figure 4: The surfaces of section (SOS; first column), and level curves of the hamiltonian (26) (second column), and Equation (34) (third column) in the cases of Aegaeon and Prometheus (top and bottom, respectively). The separatrix of the $\alpha$-regime is shown in red. Dashed-dot-horizontal green lines indicate the equilibria points obtained from hamiltonian and compared to the SOS corresponding fixed points.
4 Additional application: Aegaeon and Prometheus.
@@
Fig. 4 shows the rotational phase space around the synchronous regime in the cases of Aegaeon ($\epsilon\sim 1.5236$), and Prometheus ($\epsilon\sim 1.1727$). As pointed out in Section 1, both rotational phase spaces display the $\alpha$ and $\beta$-regimes. Inspection of the left and middle panels shows us that the position of the three equilibria points of the $\alpha$ and $\beta$-regimes obtained from level curves (middle) agree with surfaces of section in the case of Prometheus, but the same is not true in the case of Aegaeon [^7] Note that our model does not allow to obtain the loci of the separatrix of the synchronous regime since it is valid, by construction, only in the interior of the resonance..
@@
To improve the model we include terms of order higher than $J^{2}$ in the action so that the hamiltonian (26) becomes:
@@
$-\frac{1}{128\epsilon}(x^{2}+y^{2})^{3}+\frac{17}{96\sqrt{\epsilon}}ex(x^{2}+y^{2})^{2}.$ (34)
The last two terms in Equation (34) are of the order $O(J^{3})$ and $O(J^{5/2})$, respectively. The term $\frac{1}{128\epsilon}(x^{2}+y^{2})^{3}$ rises from $H_{0}^{\prime\prime}$ (the last term at the right in the Equation (8)). The second one comes from $H_{1}^{\prime\prime}$ (Equation (18)) after considering the terms of order $J^{5/2}$ at the right in Equations (10) and (11).
Inspection of Figure 4, bottom, shows a significant improvement in the case of Aaegeon. As discussed in Lei (2023), and first shown in Gkolias et al. (2019), in the specific case of the 1/1 secondary resonance, the analytical and numerical mapping of the associated fixed points diverge for values of $\epsilon>1.2$.
@@ 5 Conclusions
This work shows the results of the dynamics of rotation of out-of-round close-in small Saturnian satellites Methone and Aegaeon discovered by the Cassini spacecraft. Their shapes and physical parameters have been updated by Thomas and Helfenstein (2020). As pointed out in Section 1, the main problem we are considering in this work is related to the nature of the $\alpha$ and $\beta$-regimes of motion in the phase space of rotation of these satellites already detected for Amalthea, Metis, and Prometheus (and probably others - see Melnikov and Shevchenko 2022, and references therein).
@@
We first conduct our investigation with the well-known model of rotation developed by Goldreich and Peale (1966) and applied by many authors in distinct sorts of problems through decades (e.g. Wisdom et al. 1984, Wisdom 2004, Callegari and Rodríguez 2013). The exact equations of the 1.5 degrees of freedom model (obtained from Equation (1)) are solved numerically with the integrator provided by Everhart (1985), resulting in accurate rotational trajectories able to identify the main properties of the phase space with the surface section technique. In the case of Methone, the $\beta$-regime is located close to the origin and corresponds to the “classical” fixed point associated with the synchronous rotation motion where the amplitude of oscillation of the angle $\Psi=\theta-f$ is small at the fixed point ($\sim 1$ degree). Its center is slightly forced by an amount $\sim 0.02$ such that $\frac{\dot{\theta}}{n}\sim 1.02$, a value which can be confirmed by linear theory. The $\alpha$-regime engulfs the $\beta$-regime and it is located far from origin at $\frac{\dot{\theta}}{n}\sim 0.45$ such that $\frac{\dot{\theta}}{n}$ oscillates around 1. The amplitude of variation of $\Psi$ is large at the fixed point of the $\alpha$-regime ($\sim 30$ degree). The numerical results of the amplitude of oscillations agree with analytical estimative (Section 3.5).
@@
To understand the complexity of such rotation phase space of Methone and Aegaeon, we apply the model developed by Wisdom (2004) to study the rotation of Enceladus. Thus, we follow exactly his steps and reproduce his results; having completed this task, we could apply the methodology to our case. Equation (16) generalizes Wisdom’s theory and it is ready to write the models for 1/1 and 3/1 secondary resonances up to $J^{3}$ in the action. The terms in the expansion involving the 1/1 resonance between the frequency of physical libration $\Psi$ and the mean motion $n$ are responsible for the co-existence of the $\alpha$ and $\beta$-regimes in the rotation of the current Methone and Aegaeon’s phase spaces. Therefore, we have shown the existence of a secondary resonance within the synchronism, completing our initial goals.
@@
Figures 2 and 3 show the dependence of the domains of the $\alpha$ and $\beta$-regimes with the shape parameter of Methone. The $\beta$-regime occurs at $\epsilon=\epsilon_{c}\sim 1.0149$ such that for $\epsilon<\epsilon_{c}$ only the $\alpha$-regime exists. While there is good agreement between numerical and analytical calculations up to $J^{2}$ in the case of Methone ($\epsilon\sim 1.077$), the same is not true for Aegaeon ($\epsilon\sim 1.5236$), where the model is useful only when higher order terms are included in the hamiltonian since $\epsilon>1.2$, a superior limit for convergence of analytical estimative (Gkolias et al. 2019), as it is shown in Figure 4.
@@
As we pointed out above, trajectories within the $\alpha$-regime of motion are forced and suffer large oscillations even in the case of initial conditions located over the fixed point. Being that the current rotational states of Saturnian close-in small satellites are probably synchronous (e.g. Thomas and Helfenstein 2020), our results can be taken into account in evolutionary studies of the rotation of these bodies. Tidal models (e.g. Ferraz-Mello et al. 2008, Ferraz-Mello 2013) can be applied to estimate the final destiny of the rotation of such small bodies and also the role of the secondary resonances on its thermal emission (Wisdom 2004).
@@ ACKNOWLEDGEMENTS.
The initial results of this work has been presented at the ‘XXI Brazilian Colloquium on Orbital Dynamics’ (INPE, from December 12 to 16, 2022); and at the ‘New Frontiers of Celestial Mechanics: theory and applications’ (Department of Mathematics Tullio Levi-Civita, University of Padua, from February 15 to 17, 2023).
Appendix: Dependence of the phase space of rotation of Methone with the orbital eccentricity.
Consider two ranges of eccentricity: $A=[0.0005,0.005]$ and $B=[0.005,0.02]$.
Fig. 5 shows levels of the hamiltonian (26) for three values of eccentricity in the interval $A$. The domains of the $\alpha$ and $\beta$-regimes remain almost unchanged for eccentricities closer to the current one. For $e$ slightly larger than the current one ($e=0.005)$, the domain of the $\alpha$-regime (the secondary resonance) is enlarged, while the domain of the $\beta$-regime slightly decreases. This result is confirmed with the surface of section (SOS). Note that inspection of the two plots of bottom figures shows us that no chaotic variation is shown around the separatrix of the secondary resonance (red curve).
Discussions given above on the dependence of the $\alpha$ and $\beta$-regimes with the orbital eccentricity are restricted within the domains of the synchronous regime. To understand the role of larger orbital eccentricity in the global structure of the phase space we must consider the volume encompassing the whole synchronous resonance. Fig. 6 shows SOS for some values of eccentricity given in the interval $B$. The chaotic layer of the synchronous resonance increases as a function of the eccentricity. The domain of the $\alpha$-regime remains almost unchanged while the $\beta$-resonance is diminished. For eccentricities $e\sim 0.018$, the $\beta$-regime becomes chaotic, and for $e\sim 0.02$ it is overlapped with the synchronous chaotic layer.
Figure 5: Level curves of hamiltonian (26) for $\epsilon=1.077$ for different values of eccentricity. The plot given at right shows the surfaces of sections of solutions of hamiltonian (1) in the case $\epsilon=1.077$, $e=0.005$.
Figure 6: Surfaces of sections of solutions of hamiltonian (1) for $\epsilon=1.077$ and different values of eccentricity indicated at the top of the plots.